OpenStudy (anonymous):

determine the solution of the system y = x + 6 y = 2x

OpenStudy (anonymous):

2x=x+6, x=6, y=6+6, y=12, (6,12)

OpenStudy (anonymous):

there you go, (6,12)

OpenStudy (anonymous):

how'd you get the answer for the first one?

OpenStudy (anonymous):

well, you have to substitute either one, like you can substitute y to 2x, then after you get 2x=x+6, you subtract the x and get x=6, from there substitute to the second the x and you get y=12

OpenStudy (anonymous):

oh okay. thank you!

OpenStudy (anonymous):

no problem